If 45.0 L of natural gas, which is essentially methane (CH4), undergoes
complete combustion at 730 mm Hg and 20ºC, how many grams of each
product are formed?
1
Expert's answer
2020-12-06T13:56:22-0500

CH4 + 2 O2 = CO2 + 2 H2O


(45L CH4) (1LCO2/ 1L CH4) = 45L CO2


PV = n RT. n = (mass/ molar mass)


PV = MRT. M = (PVM/ RT)


M = ((730mmHg)(45)(44g/mol))/(62.4mmHg)(293k) =79.1 =80g CO2


(45L CH4 )((2LH2O/1LCH4 ) = 90L H2O


M = PVM/ RT = (730)(90)(18.02)/(42.4L mmHg) (293)


64.8= 65 = 70 H2O

60g H2O